OpenStudy (anonymous):

find the product (3x-8)(3x+8y)

OpenStudy (anonymous):

9x^2-24x+24xy+64y^2

OpenStudy (anonymous):

9x^2 + 24x-24x-64y

OpenStudy (anonymous):

*+24xy

OpenStudy (anonymous):

(a-b)(a+b) = a^2-b^2 (3x-8)(3x+8) = (3x)^2-8^2=9x^2-64